In 2 years, Jorge will be 3 times as old as maria is now.
In 10 years, Jorge will be 4 times as old as maria is now

Answer by mananth(16946)

let jorge's age be x
Maria's age be y
x+2=3y
x+10=4y
x+10 -x-2=4y-3y
10-2=y
y=8 Maria's age
x+2= 3*8
x=24-2
x=22 Jorge's age


3